OpenStudy (anonymous):

Solve 2x + y = 7 x - y = -4 (1, 5) (-1, 5) (1, -5) (-1, -5)

OpenStudy (anonymous):

Does the method matter?

OpenStudy (apolloschariot):

Just like the last one, you can solve this problem using substitution or elimination.

OpenStudy (anonymous):

If you use elimination, you can add them right away, and eliminate y right off the bat.

OpenStudy (anonymous):

elimination

OpenStudy (anonymous):

So any clue how to start? Look at my comment above :P

OpenStudy (anonymous):

2x + y = 7 - x - y = -4

OpenStudy (anonymous):

Add, not subtract

OpenStudy (anonymous):

If you add both equations up, you are left with 3x=3. Because the ys are cancelled out. X = 1... Then substitute into the second equation 1-y=-4. y = -5. Done.

OpenStudy (anonymous):

2x + y = 7 + x - y = -4 ____________________ 3x = 3 x = 1

OpenStudy (anonymous):

So i can eliminate answers B & D

OpenStudy (anonymous):

That's right! Now substitute x in one of the original equations, and solve for y :)

OpenStudy (anonymous):

2(1) + y = 7 2 + y = 7 -2 -2 _____________ y = 5

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

Thank you for helping!!

OpenStudy (anonymous):

Np :P